# The Dark Dimension, the Swampland, and the Dark Matter Fraction Composed of Primordial Near-Extremal Black Holes

### v1

@@ VI Conclusions
We have studied the decay rate of near-extremal black holes within the context of the dark dimension. Using generic entropy arguments we have demonstrated that Hawking evaporation of higher-dimensional near-extremal black holes proceeds at a slower rate than the corresponding Schwarzschild black holes of the same mass. Armed with our findings we have shown that if there were 5D primordial near-extremal black holes in nature, then a PBH all-dark-matter interpretation would be possible in the mass range $10^{7}\sqrt{\beta}\lesssim M/{\rm g}\lesssim 10^{21}$, where $\beta$ is a parameter that controls the difference between mass and charge of the associated near-extremal black hole.
@@ Acknowledgements
The work of L.A.A. is supported by the U.S. National Science Foundation (NSF Grant PHY-2112527). The work of D.L. is supported by the Origins Excellence Cluster and by the German-Israel-Project (DIP) on Holography and the Swampland.
@@
For completness, in this Appendix we provide of concise summary of the salient characteristics of 4D RN-dS black holes. For $d=4$, the blackening function (13) is given by
@@
Figure 1: The family of 4D RN-dS black holes. The gray shaded region represents the physical phase space of sub-extremal solutions. The boundary of this allow region has two branches: the left (or cold) branch corresponds to RN-dS extremal black holes and the right branch corresponds to charged Nariai black holes, for which the event and cosmological horizons coincide. The blue star where the two branches intersect, stands for the ultracold solution. The dashed line indicates the lukewarm solutions with ${\cal M}=|{\cal Q}|$, where the Cauchy and event horizons have the same temperature. The ${\cal Q}=0$ axis of neutral black holes indicates a big crunch singularity. Adapted from Montero:2019ekk .

### v2

@@ VI Conclusions
We have studied the decay rate of near-extremal black holes within the context of the dark dimension. Using generic entropy arguments we have demonstrated that Hawking evaporation of higher-dimensional near-extremal black holes proceeds at a slower rate than the corresponding Schwarzschild black holes of the same mass. Armed with our findings we have shown that if there were 5D primordial near-extremal black holes in nature, then a PBH all-dark-matter interpretation would be possible in the mass range $10^{5}\sqrt{\beta}\lesssim M/{\rm g}\lesssim 10^{21}$, where $\beta$ is a parameter that controls the difference between mass and charge of the associated near-extremal black hole.
@@ Acknowledgements
The work of L.A.A. is supported by the U.S. National Science Foundation (NSF Grant PHY-2112527). I.A. is supported by the Second Century Fund (C2F), Chulalongkorn University. The work of D.L. is supported by the Origins Excellence Cluster and by the German-Israel-Project (DIP) on Holography and the Swampland.
@@
For completness, in this Appendix we provide a concise summary of the salient characteristics of 4D RN-dS black holes. For $d=4$, the blackening function (13) is given by
@@
Figure 1: The family of 4D RN-dS black holes. The gray shaded region represents the physical phase space of sub-extremal solutions. The boundary of this allowed region has two branches: the left (or cold) branch corresponds to RN-dS extremal black holes and the right branch corresponds to charged Nariai black holes, for which the event and cosmological horizons coincide. The blue star where the two branches intersect, stands for the ultracold solution. The dashed line indicates the lukewarm solutions with ${\cal M}=|{\cal Q}|$, where the Cauchy and event horizons have the same temperature. The ${\cal Q}=0$ axis of neutral black holes indicates a big crunch singularity. Adapted from Montero:2019ekk .